# Fluid flow channeling and mass transport with discontinuous porosity distribution

### v3

@@ Abstract.
The flow of fluids within porous rocks is an important process with numerous applications in Earth sciences. Modeling the compaction-driven fluid flow requires the solution of coupled nonlinear partial differential equations that account for the fluid flow and the solid deformation within the porous medium. Despite the nonlinear relation of porosity and permeability that is commonly encountered, natural data show evidence of channelized fluid flow in rocks that have an overall layered structure. Layers of different rock types routinely have discontinuous hydraulic and mechanical properties. We present numerical results obtained by a novel space-time method which can handle discontinuous initial porosity (and permeability) distributions. The space-time method enables a straightforward coupling to models of mass transport for trace elements. Our results show the influence of different kinds of layering in the development of fluid-rich channels and mass transport.
@@
S.B. has been funded in part by the German Research Foundation (DFG) – project number 442047500 – SFB 1481. E.M. would like to acknowledge the German Research Foundation (DFG) – project number 521637679 for financial support. M.B. acknowledges support by the German Research Foundation (DFG) – project number 442047500 – SFB 1481.
@@ 1. Introduction
Here, we consider a poro-viscoelastic model that generalizes the one introduced in [4, 29] for the interaction of porosity and pressure. For modeling sharp transitions between materials, as caused, for example, by stacked rock layers, it is important to be able to treat porosities with jump discontinuities. These discontinuities turn out to be determined mainly by the initial condition, as it was shown in [2] based on results from [23] for smooth initial porosities. In addition, we utilize a newly-developed space-time method that has been shown to be more accurate in solving this particular problem in the presence of jump discontinuities [1]. Our approach can be used to benchmark methods that do not include discontinuities and quantify the error between the two approaches. An additional advantage of the space-time method is that it can be coupled to simple models of chemical-tracer transport [19, 11] in a straightforward manner. The results obtained from this coupling allow us to investigate the evolution of chemical anomalies in the presence of channelized fluid flow. In particular, our results are relevant to the formation of ore deposits and to the transport of trace elements in the subsurface.
@@ 1.1. The governing equations
$\partial_{t}\phi$ $=-(1-\phi)\left(\phi^{m}\frac{p}{\sigma(p)}+Q\partial_{t}p\right),$ (1a)
$\partial_{t}p$ $=\frac{1}{Q}\left(\nabla\cdot\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}\phi^{n}(\nabla p+(1-\phi)\delta\!\rho g\vec{e}_{d})-\phi^{m}\frac{p}{\eta_{\mathrm{b}}\sigma(p)}\right),$ (1b)
where $\phi$ is the porosity (void ratio), $p$ is the effective pressure, $\sigma$ accounts for decompaction weakening [16, 15], $Q$ is the compressibility (equal to $K^{-1}$, where $K$ is the bulk modulus), $\delta\!\rho=\rho^{\mathrm{s}}-\rho^{\mathrm{f}}$ the density difference, and $\frac{\eta_{\mathrm{b}}\sigma}{(1-\phi)\phi^{m}}$ can be regarded as effective viscosity similar to [27]. Finally, $t$ is time and $\vec{e}_{d}$ is the vector indicating the gravity acceleration direction (all symbols and the respective units are given in Table 1). The problem is furthermore supplemented with initial porosity $\phi_{0}\colon\Omega\to(0,1)$ and initial effective pressure $p_{0}\colon\Omega\to\mathbb{R}$.
@@
$n$ Carman-Kozeny exponent $3$
$m$ viscosity exponent $2$
@@
An extension of the hydro-mechanical model (1) is to consider the transport of a chemical tracer (such as a trace element) as described in [11, Sec. 3]. The amount of tracer is quantified using the total concentration
$\mathcal{C}:=\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}+(1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}\,,$
@@
$\partial_{t}\mathcal{C}+\nabla\cdot(\mathbf{v}^{\mathrm{e}}\mathcal{C})=0\,.$ (2)
@@
$\mathbf{v}^{\mathrm{e}}=\frac{\mathbf{v}^{\mathrm{f}}\phi}{\phi+(1-\phi)K_{D}}\,,\qquad\mathbf{v}^{\mathrm{f}}=\frac{1}{\phi}\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}\phi^{n}\big{(}\nabla p+(1-\phi)\delta\!\rho g\vec{e}_{d}\big{)}\,,$
where $K_{D}=\frac{\rho^{\mathrm{s}}\chi^{\mathrm{s}}}{\rho^{\mathrm{f}}\chi^{\mathrm{f}}}$ describes the concentration ratio of the tracer which is assumed to be constant as already indicated in Table 1. The specific form of (2) is valid at the limit where grain-scale chemical diffusion and hydro-dynamic dispersion are ignored. Previous studies indicate that, at the large scales considered here, these phenomena can be neglected [19, 26].
@@ 1.2. Existing numerical methods
Various methods have been proposed to solve the above type of problem numerically, for example finite difference schemes with implicit time-stepping in [4] and adaptive wavelets in [29]. In a number of recent works, pseudo-transient schemes based on explicit time stepping in a pseudo-time variable have been investigated. Due to their compact stencils, low communication overhead and simple implementation, such schemes are well suited for parallel computing on GPUs, so that very high grid resolutions can be achieved to compensate the low order of convergence, as shown for example in [16, 28, 17, 15, 32, 18]. Even though all of these schemes are observed to work well for smooth initial porosities $\phi_{0}$, their convergence can be very slow in problems with non-smooth $\phi_{0}$, in particular in the presence of discontinuities [1]. In such cases, due to the smoothing that is implicit in the finite difference schemes, accurately resolving sharp localized features can require extremely large grids.
@@ 1.3. Novel contributions
We solve the hydro-mechanical (HM) problem given by (1) using the space-time adaptive method introduced in [1]. This approach can handle discontinuities in the porosity $\phi$ without approximating it by a continuous function with steep gradient. As a result, smaller grids and less computational effort are needed compared to continuous schemes such as finite differences. Since the method generates efficient approximations of the entire time history of a solution to (1) in a sparse format, its coupling to the problem of chemical-tracer transport (CT) given by (2) becomes straightforward. This is because the CT problem does not give feedback to the model (1), and thus solving it can be seen as post-processing step.
@@ 2.1. Hydro-mechanical model (HM)
To solve (1) we consider the space-time adaptive method which was introduced in [1] based on a combination of a Picard iterations for (1a) and a particular adaptive least squares discretization of (1b) which itself is based on [7, 8, 9]. To make this more precise we start by introducing the new variable $\varphi=-\log(1-\phi)$, so that $\phi=1-e^{-\varphi}$, because then the system (1) can be written in the form
$\partial_{t}\varphi$ $=-\left(\beta(\varphi)\frac{p}{\sigma(p)}+Q\partial_{t}p\right),$ (3a)
$\partial_{t}p$ $=\frac{1}{Q}\left(\nabla\cdot\alpha(\varphi)(\nabla p+\zeta(\varphi))-\beta(\varphi)\frac{p}{\sigma(p)}\right),$ (3b)
@@
$\alpha(\varphi):=\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}(1-e^{-\varphi})^{n},\quad\beta(\varphi):=\frac{1}{\eta_{\mathrm{b}}}(1-e^{-\varphi})^{m},\quad\zeta(\varphi):=e^{-\varphi}\delta\!\rho g\vec{e}_{d}\,.$
@@
$\partial_{t}p^{(k)}=\frac{1}{Q}\left(\nabla\cdot\alpha(\overline{\varphi})\,(\nabla p^{(k)}+\zeta(\overline{\varphi}))-\beta(\overline{\varphi})\frac{p^{(k)}}{\sigma(p^{(k-1)})}\right),$ (4)
@@
$G[\overline{p}](p,\psi):=\begin{pmatrix}\frac{1}{Q}\big{(}\mathrm{div}(p,\psi)+\beta(\overline{\varphi})\,\frac{p}{\sigma(\overline{p})}\big{)}\\ \psi+\alpha(\overline{\varphi})\,\nabla_{\!x}p\\ p(0,\cdot)\end{pmatrix},\qquad R:=\begin{pmatrix}0\\ -\alpha(\overline{\varphi})\,\zeta(\overline{\varphi})\\ p_{0}\end{pmatrix},$
with $\mathrm{div}(p,\psi):=\partial_{t}p+\mathrm{div}_{x}\psi$ we can reformulate (4) as first-order system
@@
$\varphi^{(k+1)}(t,\cdot)=Q(p_{0}-p[\varphi^{(k)}](t,\cdot))-\int_{0}^{t}\beta(\varphi^{(k)})\frac{p[\varphi^{(k)}](s,\cdot)}{\sigma(p[\varphi^{(k)}](s,\cdot))}\,\mathrm{d}s.$
which is based on integrating (3a) in time. For more details including proofs of convergence we refer to [1, Sec. 3, 4].
The resulting adaptive scheme yields efficient approximations of localized features of solutions, in particular in the presence of discontinuities. In addition, this scheme can generate space-time grids corresponding to spatially adapted time steps; an example of this can be found in Figure 1.
Figure 1. Example of a porosity channel at $T=1.5\,\mathrm{Myr}$ (a) with the associated adaptive space-time grid (b)
The method provides computable a-posteriori estimates of the error with respect to the exact solution of the coupled nonlinear system (3). Furthermore, one obtains optimal convergence rates for $\phi$ and $p$ independent of the presence of discontinuities in $\phi$, as observed in [1, Sec. 5.2].
@@ 2.2. Chemical-tracer transport model (CT)
To solve the chemical transport equation (2), we follow its characteristics for many different starting values using an explicit Euler scheme. Note that this approach is highly parallelizable, since we need to solve a high number of independent ODEs for each starting value. By exploiting this we usually achieve very low wall-clock times, even for many starting values corresponding to a high resolution.

### v4

@@ Abstract.
The flow of fluids within porous rocks is an important process with numerous applications in Earth sciences. Modeling the compaction-driven fluid flow requires the solution of coupled nonlinear partial differential equations that account for the fluid flow and the solid deformation within the porous medium. Despite the nonlinear relation of porosity and permeability that is commonly encountered, natural data show evidence of channelized fluid flow in rocks that have an overall layered structure. Layers of different rock types have discontinuous hydraulic and mechanical properties. We present numerical results obtained by a novel space-time method, which can handle discontinuous initial porosity (and permeability) distributions efficiently. The space-time method enables straightforward coupling to models of mass transport for trace elements. Our results indicate that, under certain conditions, the discontinuity of the initial porosity influences the distribution of incompatible trace elements, leading to sharp concentration gradients and large degrees of elemental enrichment. Finally, our results indicate that the enrichment of trace elements depends not only on the channelization of the flow but also on the interaction of fluid-filled channels with layers of different porosity and permeability.
@@
The authors would like to thank the reviewers for comments that allowed us to improve the manuscript. S.B. has been funded in part by the German Research Foundation (DFG) – project number 442047500 – SFB 1481. E.M. would like to acknowledge the German Research Foundation (DFG) – project number 521637679 for financial support and Prof. Y. Podladchikov for discussions regarding the discontinuous solutions. M.B. acknowledges support by the German Research Foundation (DFG) – project number 442047500 – SFB 1481.
@@ 1. Introduction
Here, we consider a poro-viscoelastic model that generalizes the one introduced in [7, 35] for the interaction of porosity and pressure. For modeling sharp transitions between materials, as caused, for example, by stacked rock layers, it is important to be able to treat porosities with jump discontinuities. These discontinuities turn out to be determined mainly by the initial condition, as it was shown in [4] based on results from [28] for smooth initial porosities. In addition, we utilize a newly-developed space-time method that has been shown to be more accurate in solving this particular problem in the presence of jump discontinuities [3]. Our approach can be used to benchmark methods that do not include discontinuities and quantify the error between the two approaches. An additional advantage of the space-time method is that it can be coupled to simple models of chemical-tracer transport (see, for example, [23, 14]) as a post-processing step, since the entire porosity-pressure history is saved and the chemical transport problem does not feedback into the porosity-pressure (hydro-mechanical) model. The results obtained from this coupling allow us, for the first time, to investigate the evolution of chemical anomalies in the presence of channelized fluid flow, and their interactions with porosity/permeability discontinuities. In particular, our results are relevant to the formation of ore deposits and to the transport of trace elements in the subsurface.
@@ 1.1. The governing equations
$\partial_{t}\phi$ $=-(1-\phi)\left(\phi^{m}\frac{p}{\eta_{\mathrm{b}}\sigma(p)}+Q\partial_{t}p\right)\,$, $\phi(0,\cdot)=\phi_{0}\,$, (1a)
$\partial_{t}p$ $=\frac{1}{Q}\left(\mathrm{div}_{x}\bigg(\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}\phi^{n}(\nabla_{\!x}p+(1-\phi)\delta\!\rho g\vec{e}_{d})\bigg)-\phi^{m}\frac{p}{\eta_{\mathrm{b}}\sigma(p)}\right)\,$, $p(0,\cdot)=p_{0}\,$, (1b)
as previously described in [7, 35]. Here, $\phi$ denotes the porosity (void ratio), $p$ is the effective pressure, $\sigma$ accounts for decompaction weakening [20, 19], $Q$ is the compressibility (equal to $K^{-1}$, where $K$ is the bulk modulus), and $\delta\!\rho=\rho^{\mathrm{s}}-\rho^{\mathrm{f}}$ the density difference. Furthermore, $\nabla_{x}f=(\partial_{x_{1}}f,\ldots,\partial_{x_{d}}f)^{\top}$ and $\mathrm{div}_{x}\textbf{f}=\sum_{i=1}^{d}\partial_{x_{i}}f_{i}$ for functions $f:\mathbb{R}^{d}\to\mathbb{R}$, $\textbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as usual. For any function $g(t,x)$, we denote $g(0,\cdot)$ as the function $g$ at a fixed time $t=0$ with varying $x$. Finally, $t$ is time and $\vec{e}_{d}$ is the vector indicating the gravity acceleration direction (all symbols and the respective units are given in Table 1). The problem is furthermore supplemented with initial porosity $\phi_{0}\colon\Omega\to(0,1)$ and initial effective pressure $p_{0}\colon\Omega\to\mathbb{R}$.
@@
$n$ Carman-Kozeny exponent 33
$m$ viscosity exponent 22
@@
An extension of the hydro-mechanical model (1) is to consider the transport of a chemical tracer (such as a trace element) as described in [14, Sec. 3]. In particular, the trace-element transport equations are chosen since we consider that the abundance of trace elements does not affect the mechanical or the hydraulic properties of the rock. As a consequence, the trace-element transport problem depends on the hydromechanical problem, but the opposite is not true. This allows us to treat the chemical transport problem as a post-processing step after we have calculated the respective fluid velocities and the porosity distribution. The amount of tracer is quantified using the total concentration
$\mathcal{C}=\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}+(1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}\,$, (2)
@@
$\partial_{t}\mathcal{C}+\mathrm{div}_{x}(\mathbf{v}^{\mathrm{e}}\mathcal{C})=0\,.$ (3)
@@
$\mathbf{v}^{\mathrm{e}}=\frac{\mathbf{v}^{\mathrm{f}}\phi}{\phi+(1-\phi)K_{D}}\,,\qquad\mathbf{v}^{\mathrm{f}}=\frac{1}{\phi}\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}\phi^{n}\big(\nabla_{\!x}p+(1-\phi)\delta\!\rho g\vec{e}_{d}\big)\,$, (4)
where $K_{D}=\frac{\rho^{\mathrm{s}}\chi^{\mathrm{s}}}{\rho^{\mathrm{f}}\chi^{\mathrm{f}}}$ describes the concentration ratio of the tracer which is assumed to be constant as already indicated in Table 1. Note that, in this case, $K_{D}$ is a ratio of concentrations and not of mass fractions. Furthermore, equation  (3) assumes that porosity is continuous and its derivation can be found in Appendix A. For cases where porosity is discontinuous, the jump condition must guarantee the conservation of mass at the discontinuity (see Appendix B for details).
1.2. Applicability of assumptions
The previous hydro-mechanical system of equations  (1) results from the simplification of the multiphase-, viscoelastic-Stokes’ equations at the static limit. The static limit occurs when no far-field stresses are imposed at the boundaries, and the buoyancy stresses are relatively small within the model domain. This limit is justified in cases where the effective pressure is close to zero. In such cases, the shear stresses that rocks can support are very small and, in many applications, can be assumed to be negligible [1, 7, 8, 27]. Being close to the static limit implies that the solid velocity for the mechanical problem is taken at the limit where $\mathbf{v}^{\mathrm{f}}\gg\mathbf{v}^{\mathrm{s}}\approx 0$ (but generally $\mathrm{div}_{x}\mathbf{v}^{\mathrm{s}}\neq 0$).
One particular process where the trace-element transport is important is when melt is ascending within the Earth’s mantle [23]. In regions such as in the mantle wedge or within a mantle plume, the temperature does not change significantly. In these geodynamic environments, the confining pressure is large and the melt-filled porosity of the mantle rock is very small, typically in the order of 0.001-0.01 [29]. To model the trace-element equilibrium and the chemical interaction between solid and fluid, we use the partition coefficient $K_{D}$. The partition coefficient changes as a function of the mineralogy of the rock, its pressure and its temperature. However, for a given material, the variation of the partition coefficient with pressure is very weak and can be considered constant over several GPa of pressure [33].
Having $\mathbf{v}^{\mathrm{f}}$ from (1) allows the solution of (3). The specific form of (3) is valid at the limit where grain-scale chemical diffusion and hydro-dynamic dispersion are ignored. Previous studies indicate that, on the large scales considered here, these phenomena can be neglected [23, 32].
@@ 1.3. Existing numerical methods
Various methods have been proposed to solve the hydro-mechanical problem (1) numerically, for example finite difference schemes with implicit time-stepping in [7] and adaptive wavelets in [35]. In a number of recent works, pseudo-transient schemes based on explicit time stepping in a pseudo-time variable have been investigated. Due to their compact stencils, low communication overhead and simple implementation, such schemes are well suited for parallel computing on GPUs, so that very high grid resolutions can be achieved to compensate the low order of convergence, as shown for example in [20, 34, 21, 19, 38, 22]. Even though all of these schemes are observed to work well for smooth initial porosities $\phi_{0}$, their convergence can be very slow in problems with non-smooth $\phi_{0}$, in particular in the presence of discontinuities [3]. Examples of this behavior are also shown in Appendix C. In such cases, due to the smoothing that is implicit in the finite difference schemes, accurately resolving sharp localized features can require extremely large grids that are computationally inefficient.
@@ 1.4. Novel contributions
Our approach considers the utilization of a space-time method to solve the hydro-mechanical problem (1). This method was introduced for this particular problem in  [3], and has the advantage that the entire solutions of porosity and effective pressure fields can be stored in a space-time grid. In addition, this approach can handle discontinuities in the porosity $\phi$ without approximating it by a continuous function with steep gradient. As a result, smaller grids and less computational effort are needed compared to continuous schemes such as finite differences. Since the method generates efficient approximations of the entire time history of a solution to (1) in a sparse format, its coupling to the problem of chemical-tracer transport (CT) given by (3) becomes straightforward. This is because the CT problem does not give feedback to the model (1), and thus solving it can be seen as post-processing step.
@@ 2.1. Hydro-mechanical model (HM)
To solve (1) we consider the space-time adaptive method which was introduced in [3] based on a combination of a Picard iterations for (1a) and a particular adaptive least squares discretization of (1b) which itself is based on [10, 11, 12]. To make this more precise, we start by introducing the new variable
$\varphi=-\log(1-\phi)\,$, (5)
so that $\phi=1-e^{-\varphi}$. The previous transformation allows the investigation of cases where the porosity is larger than the typical ”small-porosity limit”[35]. The system (1) can then be written in the form
$\partial_{t}\varphi$ $=-\left(\beta(\varphi)\frac{p}{\sigma(p)}+Q\partial_{t}p\right)\,$, $\varphi(0,\cdot)=\varphi_{0}\,$, (6a)
$\partial_{t}p$ $=\frac{1}{Q}\left(\mathrm{div}_{x}\Big(\alpha(\varphi)(\nabla_{\!x}p+\zeta(\varphi))\Big)-\beta(\varphi)\frac{p}{\sigma(p)}\right)\,$, $p(0,\cdot)=p_{0}\,$, (6b)
@@
$\alpha(\varphi)=\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}(1-e^{-\varphi})^{n},\quad\beta(\varphi)=\frac{1}{\eta_{\mathrm{b}}}(1-e^{-\varphi})^{m},\quad\zeta(\varphi)=e^{-\varphi}\delta\!\rho g\vec{e}_{d}\,.$ (7)
@@
$\partial_{t}p^{(k)}=\frac{1}{Q}\left(\mathrm{div}_{x}\Big(\alpha(\overline{\varphi})\,(\nabla_{\!x}p^{(k)}+\zeta(\overline{\varphi}))\Big)-\beta(\overline{\varphi})\frac{p^{(k)}}{\sigma(p^{(k-1)})}\right)\,,\quad p^{(k)}(0,\cdot)=p_{0}\,$, (8)
@@
$G[p^{(k-1)}](p^{(k)},\psi^{(k)})=\begin{pmatrix}\frac{1}{Q}\Big(\mathrm{div}(p^{(k)},\psi^{(k)})+\beta(\overline{\varphi})\,\frac{p^{(k)}}{\sigma(p^{(k-1)})}\Big)\\ \psi^{(k)}+\alpha(\overline{\varphi})\,\nabla_{\!x}p^{(k)}\\ p^{(k)}(0,\cdot)\end{pmatrix},\quad R=\begin{pmatrix}0\\ -\alpha(\overline{\varphi})\,\zeta(\overline{\varphi})\\ p_{0}\end{pmatrix}$, (9)
with $\mathrm{div}(p,\psi)=\partial_{t}p+\mathrm{div}_{x}\psi$, we can reformulate (8) as first-order system
@@
Note that the second row of (9) can be rewritten as $\psi^{(k)}=-\alpha(\overline{\varphi})(\nabla_{\!x}p^{(k)}+\zeta(\overline{\varphi}))$. Plugging $\psi^{(k)}$ into $\mathrm{div}(p^{(k)},\psi^{(k)})$ in the first row of (9) yields (8).
@@
$\varphi^{(k+1)}(t,\cdot)=Q(p_{0}-p[\varphi^{(k)}](t,\cdot))-\int_{0}^{t}\beta(\varphi^{(k)})\frac{p[\varphi^{(k)}](s,\cdot)}{\sigma(p[\varphi^{(k)}](s,\cdot))}\,\mathrm{d}s.$ (11)
which is based on integrating (6a) in time. For more details including proofs of convergence we refer to [3, Sec. 3, 4].
This scheme can generate space-time grids corresponding to spatially adapted time steps; an example of this can be found in Figure 1.
Figure 1. Example of a porosity channel at $T=1.5\,\mathrm{Myr}$ (a) with the associated adaptive space-time grid (b); the color of each grid cell denotes its refinement level
Furthermore, the method provides computable a-posteriori estimates of the error with respect to the exact solution of the coupled nonlinear system (6). Therefore, this method can be used to steer an adaptive grid refinement routine which yields efficient approximations of localized features of solutions, in particular in the presence of discontinuities. In addition, one obtains optimal convergence rates for $\phi$ and $p$ independent of the presence of discontinuities in $\phi$, as observed in [3, Sec. 5.2].
@@ 2.2. Chemical-tracer transport model (CT)
To solve the chemical transport equation (3), we follow its characteristics. Namely, we consider
$\begin{split}\partial_{t}x(t)&=\mathbf{v}^{\mathrm{e}}(t,x(t))\,,\\ c(t)&=\mathcal{C}(t,x(t))\,.\end{split}$ (12)
Then we calculate
[…]
the previous yields a coupled system of ordinary differential equations (ODEs)
[…]
that is used to solve (3). The solution is provided along the characteristics given by $\mathbf{v}^{\mathrm{e}}$ starting with some initial value $x_{0}$ and initial concentration $c_{0}$. We use an explicit Euler scheme to solve (14) for many different starting values $x_{0}$. Note that this approach is highly parallelizable, since we need to solve a high number of independent ODEs for each starting value. By exploiting this we usually achieve very low wall-clock times, even for many starting values corresponding to a high resolution. Note furthermore that this approach only conserves the quantity $\mathcal{C}$ if $\mathbf{v}^{\mathrm{e}}$ is continuous. For the discontinuous cases one needs to ensure continuity of the flux and we refer to Appendix B for more details.